Solve. 3x – 0.11 = 2(x – 0.66)
A) –1.07 B) –1.21 C) –1.42 D) –0.87

I did not come with an answer for this can you help me?

Found 2 solutions by tutorcecilia, prabhjyot:
Answer by tutorcecilia(2152)

3x – 0.11 = 2(x – 0.66)
3x-0.11=2x-1.32
3x-2x-0.11=2x-2x-1.32
x-0.11=-1.32
x-0.11+0.11=-1.32+0.11
x=-1.21

Answer by prabhjyot(164)

Solve. 3x – 0.11 = 2(x – 0.66)
A) –1.07 B) –1.21 C) –1.42 D) –0.87
Solution: 3x – 0.11 = 2(x – 0.66)
3x-0.11 = 2x- 1.32
Taking like terms to one side of the equation
3x-2x= -1.32 + 0.11
x=-1.21
ANSWER: B)-1.21